OpenStudy (anonymous):

Which literal equation is equivalent to a equals negative five plus four times b? b equals the sum of a plus five all over four b equals the a minus five all over four b equals negative four times a plus five b equals four times a plus five

OpenStudy (anonymous):

FAN AND MEDAL :>

OpenStudy (anonymous):

given; a = -5+4b from this we can find the value of b \[\frac{a+5}{4}=b\] so, b equals sum of a and 5 all over four

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You are welcome :)